At a scientific congress, 85% of participants speak english (A), 35% speak french (F) and 25% speak both of these languages. If we randomly chose a congressman, what would be the probability of him being able to speak English or French?
Answer by ikleyn(52778)

.

Use the general formula of the probability theory


    P(A U B) = P(A) + P(B) - P(A and B).


In your case it takes the form


    P(English or French) = P(English) + P(French) - P(English and French) = 0.85 + 0.35 - 0.25 = 0.95 = 95%    ANSWER

Solved, answered, explained and completed.